An isosceles triangle has two angles of equal measure. In isosceles triangle ABC, angles A and B are equal in measure and each of these angles is 20 degrees more than three times the measure of angle C. Find the measure of each angle.
Answer by mananth(16946)

angle c --x degrees
angle A = 3x+20
Angle B = 3x+20
3x+20+3x+20+x=180 ( sum of nagles of triangle = 180 degrees)
7x+40=180
7x= 180-40
7x=140
/20
x= 20
Angle C = 20
Angle B = 3x+20 = 60+20=80
Angle A= 80